OpenStudy (anonymous):

Solve 2 sin x - 1 = 0 for 0 degree less than or equal to x less than 2 pi

OpenStudy (anonymous):

oh my mad that will be 2 cos^2

OpenStudy (anonymous):

not 2 sin

OpenStudy (anonymous):

sinx=1/2 sinx=sinpi/6 x=pi/6 or 5pi/6

OpenStudy (anonymous):

 which one is sin(x)=1/2